Benzene:

Kow: 1.99

Aqueous Solubility: 2000 mg/L

Vapor Pressure: 75 mmHg

Spill description: A truck containing 5000 grams of your assigned chemical has spilled into a closed lake with the following dimensions: 100 meters long, 100 meters wide, and 1 meter deep.

  1. What is the concentration of your chemical (in grams per liter) in the lake assuming it quickly reaches equilibrium and lake is filled to the brim with water? This is not something we covered specifically in class; however, you should be able to calculate a volume of the water in the lake based on the details given above. Once you have a volume you can calculate a concentration (5 points, show your calculations).
  1. Based on the aqueous solubility of your chemical, are you likely to see precipitation of your assigned chemical at this concentration? Why or why not? (5 points, 1-2 sentences)
    1. Is the season during which the spill take place likely to influence the dissolved concentration of your chemical? Why? (5 points, 1-2 sentences)
      1. The sediments in this lake are known to have high levels of organic carbon. Provide a justification of whether this will influence the fate of your chemical based on the characteristics above (5 points, 1-2 sentences).
        1. Assuming that some portion of your chemical will partition to sediment, what are the factors that can result in release of the chemical from the sediment (7.5 points, 2-3 sentences)?

                        

          1. What are the likely routes of chemical degradation in the water? Is the turbidity of the water likely to influence your chemicals degradation? (7.5 points, 2-3 sentences)
            1. Based on what you know about chemical fate in air and vapor pressure, is any of your chemical likely to partition to the air? Please provide an explanation of why or why not? (5 points, 1-2 sentences)
              1. Based on the characteristics listed above, where would you expect the primary fate of each chemical to be: Air, water, or sediment. Please provide an explanation of your reasoning. (10 points, 3-4 sentences)
              Answer #1

              What is the concentration of your chemical (in grams per liter) in the lake assuming it quickly reaches equilibrium and lake is filled to the brim with water?

              Volume of the water in the lake = 100 meters * 100 meters * 1 meter = 104 meter3 = 104 * 103 Liters = 107 Liters

              [ 1 meter3 = 103 L]

              Concentration of the chemical is the lake water = Mass of the chemical/ Volume of water = 5000 g / 107 L = 5000 * 103 mg / 107 L = 0.5 mg / L

              Based on the aqueous solubility of your chemical, are you likely to see precipitation of your assigned chemical at this concentration? Why or why not?

              The aqueous solubility of the chemical is 2000 mg / L. But the concentration of the chemical in lake water is only 0.5 mg / L which does not exceed the aqueous solubility. If the concentration exceeds aqueous solubility, then there occurs precipitation.

              So, here no precipitation occurs. All chemical is dissolved in lake water.

              Is the season during which the spill take place likely to influence the dissolved concentration of your chemical? Why?

              No, because all benzene are dissolved in lake water. If some would remain undissolved then the season would affect to dissolve the undissolved chemical. But as all benzene is dissolved, it is not influenced further by season.

              The sediments in this lake are known to have high levels of organic carbon. Provide a justification of whether this will influence the fate of your chemical based on the characteristics above.

              If the organic compounds in the lake water reacts with benzene, then the benzene concentration will decrease.